# From Seedling to Harvest: The GrowingSoy Dataset for Weed Detection in Soy Crops via Instance Segmentation

### v1

@@ Abstract
Deep learning, particularly Convolutional Neural Networks (CNNs), has gained significant attention for its effectiveness in computer vision, especially in agricultural tasks. Recent advancements in instance segmentation have improved image classification accuracy. In this work, we introduce a comprehensive dataset for training neural networks to detect weeds and soy plants through instance segmentation. Our dataset covers various stages of soy growth, offering a chronological perspective on weed invasion’s impact, with 1,000 meticulously annotated images. We also provide 6 state of the art models, trained in this dataset, that can understand and detect soy and weed in every stage of the plantation process. By using this dataset for weed and soy segmentation, we achieved a segmentation average precision of 79.1% and an average recall of 69.2% across all plant classes, with the YOLOv8X model. Moreover, the YOLOv8M model attained 78.7% mean average precision (mAp-50) in caruru weed segmentation, 69.7% in grassy weed segmentation, and 90.1% in soy plant segmentation. [^1] Video of the experiments is available at: https://youtu.be/Ir0vwtAk8Tc [^2] GitHub Repository is available at: https://github.com/raulsteinmetz/soy-segmentation [^3] SoyGrowing dataset is available at: https://github.com/raulsteinmetz/soy-segmentation-ds
@@ I Introduction
Effective crop management is essential for maximizing yields, especially in environments prone to issues like weed infestations, such as soybean crops. Despite efforts to control weeds through methods like chemical applications, the constant threat of unexpected weed invasions looms large, endangering crop quality and yield. Thus, vigilance and responsive agricultural practices remain paramount.
@@
Soybeans occupy a pivotal role on the world stage, being the second most produced vegetable oil and serving as a major source of high-quality protein for both animal feed and vegetable oil production, while also delivering essential protein for human consumption [1]. It is of uttermost importance to use management techniques that allow to identify and remove weeds as early as possible, increasing crop production.
@@
In recent years, rapid advancements in hardware and software technologies have propelled neural networks into the spotlight as powerful tools capable of automating crucial agricultural processes [2, 3]. These innovations encompass autonomous robot control [4, 5] and sophisticated computer vision algorithms, such as Convolutional Neural Networks (CNN) for image classification [6]. These technologies enable automated disaster detection while simultaneously conducting a comprehensive analysis of plantation status and key metrics [7].
@@
One groundbreaking development in computer vision is instance semantic segmentation [8, 9], a technique that partitions unfamiliar images into distinct components and entities with pixel-level precision, revolutionizing object recognition and delineation. This advancement holds immense promise for agricultural computer vision, where precise identification and understanding of individual components within a crop field are crucial for effective crop management. However, the lack of data properly labeled and processed for segmentation is yet a bottleneck for the development of advanced applied solutions.
@@
The prevailing issue in crop datasets lies in the lack of segmentation labels, as the majority remain unlabeled for this purpose. Notably, datasets such as DeepWeeds [10], Weed25 [11], and the Soybean images dataset  [12] are exclusively annotated for image classification, overlooking the crucial segmentation aspect. Furthermore, a significant portion of these datasets consists of images with low resolution and inadequate illumination.
@@
In contrast, our dataset address these limitations by providing a substantial number of high-quality images depicting soy and weeds. What sets our dataset apart is the inclusion of segmentation labels for all images, enhancing their uniqueness. This not only contributes to a more comprehensive dataset but also fills a crucial gap in the availability of well-annotated crop datasets for segmentation tasks.
Hence, this research aims to contribute with the context presented through the following contributions:
@@
  • We present a novel dataset with 1,000 meticulously annotated and high-resolution images for training instance segmentation neural networks, focusing on weed-infested soy plantations. The dataset spans the entire soy growth process, from initial stages to maturity with pervasive weed presence.
@@
  • Additionally, we provide six state-of-the-art neural network models and perform an in-depth comparison among them, leveraging this temporal dataset to achieve impressive mean average precision in weed and soy segmentation across all soy plantation stages.
The YOLOv5 [13] and YOLOv8 [14] models trained using our dataset, achieved promising performance. Notably, the YOLOv8m model achieved a segmentation mean average precision of 78.9% for caruru weed, 62.8% for grassy weed, and 87.7% for soy plants. The YOLOv8m model also attained segmentation map-50 scores of 78.7% for caruru weed, 69.6% for grassy weed, and 90.1% for soy plants. Additionally, the YOLOv8x model attained 0.791 precision score between all classes and 0.692 recall score between all classes. The instance segmentation precision of the model trained using our dataset is depicted in Fig. 1.
@@
The structure of this paper is as follows: Section II offers an extensive literature review, delving into previous research in computer vision for crop monitoring and instance segmentation. In Section III, we outline our methodology for creating the dataset. In Section IV, we depict the process of selecting and training the neural networks involved in the benchmark study. Section V presents the results obtained using the neural networks on our dataset. Section VI presents an analysis of the performance in images that contain both soy and weed, in seedling and harvest plantation stages. Finally, Section VII summarizes our conclusions drawn from this study, highlighting the significance of our findings for agricultural management and the potential for further advancements in instance segmentation for crop monitoring.
@@ II Related Work
In the domain of crop monitoring, another significant application of CNNs is in the detection of plant diseases. Medhi and Deb [22] introduced an image dataset featuring various banana plant varieties and the diseases associated with them, summing up to an extensive collection of over 8000 images for deep neural network training. In a similar vein, Moupojou et al. [23] introduced the FieldPlant dataset, consisting of 5,170 plant disease images acquired directly from plantations. These images were meticulously annotated at the leaf level by plant pathologists. The study demonstrates that FieldPlant surpasses current datasets in terms of performance in classification tasks, establishing itself as a more precise tool for disease detection in practical agricultural scenarios.
@@ III-B Dataset Creation and Annotation
To curate a representative database of soy plantation stages, we systematically selected 1000 frames from recorded videos. This sequential extraction, with a fixed step across all videos, ensured comprehensive coverage of each stage in the soy plantation lifecycle. Through iterative experimentation, we determined the optimal number of images for model training, ranging from 250 to 1500 images. Interestingly, discernible model improvement plateaued after reaching 1000 images, affirming that this quantity sufficiently encapsulates the task’s complexity.
Subsequently, the selected frames underwent resizing to 640x640 pixels using cubic interpolation, aligning them with the specifications required for neural network models. To provide a glimpse into the dataset, a subset of these images is showcased in Fig. 4. This meticulous curation process not only ensures diversity in representation but also optimizes model performance for the designated task.
@@
IV Neural Network Models
In this section, we delve into an analytical exploration of two paramount frameworks in the realm of computer vision segmentation: YOLOv5 [13] and YOLOv8 [14], these architectures and their different size models (medium, large and x-large) were the ones chosen for our instance segmentation approach. These frameworks, underpinned by cutting-edge research and technological advancements, have significantly shaped the trajectory of real-time object detection and segmentation methodologies.
YOLOv5 is a significant iteration in the YOLO series, known for its real-time object detection capabilities. Developed by Ultralytics, YOLOv5 became a defining model in the computer vision industry. It operates by dividing an image into a grid and predicting bounding boxes and class probabilities for each grid cell. YOLOv5 uses anchor boxes, which are predefined shapes, to predict object shapes in an image. The model’s success is attributed to its flexible Pythonic structure, which allowed the community to introduce new modeling improvements and share them quickly.
YOLOv8 is the latest state-of-the-art model in the YOLO series, designed for object detection, image classification, and instance segmentation tasks. Developed by Ultralytics, the same team behind YOLOv5, YOLOv8 introduces numerous architectural and developer experience enhancements over its predecessor. One of the significant features of YOLOv8 is its anchor-free detection, which predicts the center of an object directly instead of using an offset from a known anchor box. This approach reduces the number of box predictions, speeding up the Non-Maximum Suppression (NMS) process. Additionally, YOLOv8 has introduced new convolutions, changing the stem’s first 6x6 conv to a 3x3 and introducing the C2f module. The model also incorporates online augmentations during training, such as mosaic augmentation.
The primary distinction between YOLOv5 and YOLOv8 is their approach to object detection. While YOLOv5 relies on anchor boxes, YOLOv8 operates as an anchor-free model, predicting object centers directly. This change addresses the challenges of anchor boxes, which might not represent the distribution of a custom dataset. Furthermore, YOLOv8 has introduced new convolutions and modules, such as the C2f module, which differentiates it from YOLOv5. Another notable difference is the training routine. YOLOv8 employs online augmentations, like mosaic augmentation, which stitches four images together, enhancing the model’s ability to recognize objects in various locations and conditions. In terms of accuracy, YOLOv8 has shown improvements over YOLOv5, especially when evaluated against benchmarks like COCO and Roboflow 100.
@@ V Evaluation and Results
Figure 5: Performance metrics for YOLOv8 and YOLOv5 during training and validation over several epochs. The legends for each graph include variations of the model: ‘m’, ‘l’, and ‘x’, representing different model sizes or configurations, with ‘train’ indicating training data and ‘val’ indicating validation data.
The neural networks trained on the temporal dataset demonstrated significant precision, recall, and mean average pixel precision throughout all stages of the plantation. This section provides an overview of the network training process, presents the achieved metrics, and conducts a comparative analysis between the YOLO architectures.
@@
The learning curves can be seen in Fig. 5. In Figures 5(b) and 5(e), the class loss curves for both the validation and training sets across all models are displayed. In Figures 5(c) and 5(f), the loss curve for segmentation can be seen. The loss curve for class is measured using Distribution Focal Loss (the focal loss helps to address class imbalance in object detection by assigning different weights to easy and hard examples), and for segmentation is Binary Cross Entropy with Logits (binary classification loss function that combines sigmoid transformation of logits with binary cross-entropy calculation, it’s used for stable and efficient training in binary classification tasks). These curves display a gradual and consistent decline, highlighting the models’ efficacy in comprehending the data throughout the training process. Additionally, the mean average precision on both the training and validation sets is presented in Figures 5(a) and 5(d), showcasing a steady and continuous upward trend, ultimately converging to yield outstanding results. The mAp-50 is calculated by averaging the precision values at the top 50 ranked predictions for a given task, providing a powerful performance metric for object detection.
@@
In Table I, mAP-50 metrics for bounding boxes in the test set are shown. The YOLOv8m model achieved the highest scores, with 0.789 for caruru weed objects, 0.628 for grassy weed objects, and 0.887 for Soy objects. Notably, the YOLOv8 models outperformed the YOLOv5 models across the vast majority classes, indicating a significant improvement.
Table II presents the average bounding box precision and recall metrics for all classes across all models. YOLOv8x exhibited the highest average precision at 0.791, while YOLOv8m achieved the best recall at 0.733. Notably, YOLOv8 and YOLOv5 models demonstrated similar results in these metrics.
@@
The segmentation mAP-50 values can be seen in Table III, with YOLOv8m consistently outperforming other models across all classes, mirroring the box mAP-50 results. Specifically, YOLOv8m achieved a mAP of 0.787 for caruru and 0.696 for grassy weed, and an impressive 0.901 for Soy. Once more, the YOLOv8 models demonstrated superior performance compared to the YOLOv5 models.
@@ VI Analysis of performance on initial and late stages
In this section, the primary focus revolves around the neural network’s pivotal role in effectively detecting weed presence. Our analysis will be dedicated to evaluating the performance of the YOLOv8 models on images featuring a coexistence of both weed and soy elements, through box predictions rather then mask predictions. We meticulously curated two distinct sets of 25 images, each depicting varying degrees of weed presence. The initial group comprises images captured during the seedling and early stages of plantation, while the second group encapsulates images from the harvest stages. Our objective is to scrutinize and comprehend the YOLOv8 models’ efficacy in accurately identifying weed instances in these diverse agricultural scenarios, and if there is a several difference in performance from early to late stages.
@@
Considering the box mAP50 results outlined in Table V and Table VI, a noteworthy observation emerges. Despite YOLOv8x being notably surpassed by the more compact YOLOv8m model, a distinctive trend becomes evident in scenarios featuring both weed and soy elements, irrespective of whether the images capture the seedling or harvest stages. Notably, YOLOv8x exhibits a consistent trend of outperforming its counterparts.
@@
An additional intriguing facet, gleaned from the tabulated data, is the absence of significant disparities in model performance across distinct stages of plantation. This observation underscores the diversity of the dataset, affirming that the trained models possess the capability to comprehend and effectively navigate through a spectrum of agricultural scenarios.
Taking into consideration the box precision and recall metrics showcased in Table VII and Table VIII, it becomes apparent that no single model distinctly outperforms the others. All YOLOv8 models consistently deliver commendable results in accurately detecting both soy and weed in the images. Notably, a subtle trend emerges where precision and recall metrics exhibit a slight improvement in the later stages of plantation, although the magnitude of this enhancement is not particularly significant.
@@
In conclusion, this analysis not only underscores the similarity in performance of the networks across different stages of plantation, highlighting the dataset’s inherent diversity and the models’ robust generalization capabilities, but it also reveals a significant performance boost in images featuring the simultaneous presence of both soy and weed. This enhancement aligns seamlessly with the primary objective of the network, which is the segmentation and detection of soy and weed, reaffirming the effectiveness of the models in fulfilling their intended purpose.
VII Conclusion

### v2

@@ Abstract
Deep learning, particularly Convolutional Neural Networks (CNNs), has gained significant attention for its effectiveness in computer vision, especially in agricultural tasks. Recent advancements in instance segmentation have improved image classification accuracy. In this work, we introduce a comprehensive dataset for training neural networks to detect weeds and soy plants through instance segmentation. Our dataset covers various stages of soy growth, offering a chronological perspective on weed invasion’s impact, with 1,000 annotated images. To validate our data, we also provide 6 state of the art models, trained in this dataset, that can understand and detect soy and weed in every stage of the plantation process, the best results achieved were a segmentation average precision of 79.1% and an average recall of 73.3% across all plant classes. Moreover, the YOLOv8M model attained 78.7% mean average precision (mAp-50) in caruru weed segmentation, 69.6% in grassy weed segmentation, and 90.1% in soy plant segmentation. [^1] Video of the experiments is available at: https://youtu.be/Ir0vwtAk8Tc [^2] GitHub Repository is available at: https://github.com/raulsteinmetz/soy-segmentation [^3] SoyGrowing dataset is available at: https://github.com/raulsteinmetz/soy-segmentation-ds
@@ I Introduction
Effective crop management is essential for maximizing yields, particularly in environments susceptible to issues like weed infestations, such as soybean crops. Despite efforts to control weeds through methods such as chemical applications, the persistent threat of unexpected weed invasions remains a significant risk to crop quality and yield. Therefore, vigilant and responsive agricultural practices are crucial.
@@
Soybeans occupy a pivotal role globally, being the second most produced vegetable oil and serving as a major source of protein for both animal feed and human consumption [1]. Effective management techniques that enable early identification and removal of weeds are essential for increasing crop production.
@@
In recent years, advancements in hardware and software technologies have propelled neural networks into the spotlight as powerful tools capable of automating crucial agricultural processes [2, 3]. These innovations include autonomous robot control [4, 5] and computer vision algorithms, such as Convolutional Neural Networks (CNNs) for image classification [6]. These technologies facilitate automated disaster detection while simultaneously conducting a comprehensive analysis of plantation status and key metrics [7].
@@
A recent advancement in computer vision is instance semantic segmentation [8, 9], a method that divides images into distinct segments and entities with pixel-level accuracy, enhancing object recognition and delineation. This technique has significant potential in agricultural computer vision, where accurate identification and analysis of individual components within a crop field are essential for efficient crop management. However, the scarcity of properly labeled and processed data for segmentation remains a challenge for the development of advanced applied solutions.
@@
The prevailing issue in crop datasets lies in the lack of segmentation labels, as the majority remain unlabeled for this purpose. Notably, datasets such as DeepWeeds [10], Weed25 [11], and the Soybean images dataset  [12] are exclusively annotated for image classification, overlooking the segmentation aspect. Furthermore, a significant portion of these datasets consists of images with low resolution and inadequate illumination.
@@
In contrast, we introduce a dataset that addresses these limitations by providing a 1000 high-quality images depicting soy and weeds. What distinguishes our dataset is the inclusion of segmentation labels for all images. This enhancement contributes to a more comprehensive dataset and fills a crucial gap in the availability of well-annotated crop datasets for segmentation tasks. A set of segmented images using a model trained on our dataset is depicted in Fig. 1.
Hence, the contributions of this work are:
@@
  • We present a novel dataset with 1,000 annotated and high-resolution images for training instance segmentation neural networks, focusing on weed-infested soy plantations. The dataset spans the entire soy growth process, from initial stages to maturity with pervasive weed presence.
@@
  • Additionally, to validate the dataset, we provide six state-of-the-art neural network models and perform an in-depth comparison among them, leveraging this temporal dataset to achieve impressive mean average precision in weed and soy segmentation across all soy plantation stages.
@@
The structure of this paper is as follows: Section II offers a literature review, delving into previous research in computer vision for crop monitoring and instance segmentation. In Section III, we outline our methodology for creating the dataset. In Section IV, we depict the process of selecting and training the neural networks involved in the study. Section V presents the results obtained using the neural networks on our dataset. Section VI presents an analysis of the performance in images that contain both soy and weed, in seedling and harvest plantation stages. Section VII present some clarifications about the results and analysis. Finally, Section VIII summarizes our conclusions drawn from this study, highlighting the significance of our findings for agricultural management and the potential for further advancements in instance segmentation for crop monitoring.
@@ II Related Work
In the domain of crop monitoring, another significant application of CNNs is in the detection of plant diseases. Medhi and Deb [20] introduced an image dataset featuring various banana plant varieties and the diseases associated with them, summing up to an extensive collection of over 8000 images for deep neural network training. In a similar vein, Moupojou et al. [21] introduced the FieldPlant dataset, consisting of 5,170 plant disease images acquired directly from plantations. These images were annotated at the leaf level by plant pathologists. The study demonstrates that FieldPlant surpasses current datasets in terms of performance in classification tasks, establishing itself as a more precise tool for disease detection in practical agricultural scenarios.
@@ III-B Dataset Creation and Annotation
To curate a representative database of soy plantation stages, we selected 1000 frames from recorded videos. A sequential extraction, with a fixed step across all videos, ensured comprehensive coverage of each stage in the soy plantation life cycle. Through iterative experimentation, we determined the optimal number of images for model training, ranging from 250 to 1500 images. Interestingly, discernible model improvement plateaued after reaching 1000 images, affirming that this quantity sufficiently encapsulates the task’s complexity.
Subsequently, the selected frames, initially with 4992 x 2496 pixels, underwent resizing to 640x640 pixels using cubic interpolation, aligning them with the specifications required for neural network models. To provide a glimpse into the dataset, a subset of these images is showcased in Fig. 4.
@@
IV Models Trained for Data Validation
In this section, we analyze two significant frameworks in image segmentation, used to validate the quality of our data: YOLOv5 [28] and YOLOv8 [29]. These architectures and their various size models (medium, large, and x-large) were chosen for our instance segmentation approach.
YOLOv5, developed by Ultralytics, is recognized for its real-time object detection capabilities. It divides an image into a grid and predicts bounding boxes and class probabilities for each grid cell using anchor boxes, which are predefined shapes for predicting object shapes. YOLOv5’s success is largely due to its flexible Pythonic structure, enabling rapid community-driven improvements.
YOLOv8, the latest model in the YOLO series, also developed by Ultralytics, supports object detection, image classification, and instance segmentation. It introduces architectural enhancements over YOLOv5, including anchor-free detection that predicts object centers directly, reducing the number of box predictions and speeding up the Non-Maximum Suppression (NMS) process. YOLOv8 also features new convolutions and the C2f module, and it incorporates online augmentations during training, such as mosaic augmentation.
The primary distinction between YOLOv5 and YOLOv8 lies in their approach to object detection. YOLOv5 uses anchor boxes, while YOLOv8 employs anchor-free detection, predicting object centers directly. This approach addresses the limitations of anchor boxes, which may not represent the distribution of custom datasets. YOLOv8 also introduces new convolutions and modules, such as the C2f module, and utilizes online augmentations like mosaic augmentation, enhancing the model’s ability to recognize objects under various conditions. In terms of accuracy, YOLOv8 has shown improvements over YOLOv5, particularly on benchmarks like COCO and Roboflow 100.
@@ V Evaluation and Results
Figure 5: Performance metrics for the YOLOv8 and YOLOv5 models during training and validation.
The neural networks trained on the temporal dataset demonstrated significant precision, recall, and mean average pixel precision throughout all stages of the plantation. This section provides an overview of the network training process and presents the achieved metrics.
@@
The learning curves can be seen in Fig. 5. In Figures 5(b) and 5(e), the class loss curves for both the validation and training sets across all models are displayed. In Figures 5(c) and 5(f), the loss curve for segmentation can be seen. The loss curve for class is measured using Distribution Focal Loss (the focal loss helps to address class imbalance in object detection by assigning different weights to easy and hard examples), and for segmentation is Binary Cross Entropy with Logits (binary classification loss function that combines sigmoid transformation of logits with binary cross-entropy calculation, it’s used for stable and efficient training in binary classification tasks). These curves display a gradual and consistent decline, highlighting the models’ efficacy in comprehending the data throughout the training process. Additionally, the mean average precision on both the training and validation sets is presented in Figures 5(a) and 5(d), showcasing a steady and continuous upward trend, ultimately converging to yield strong results. The mAp-50 is calculated by averaging the precision values at the top 50 ranked predictions for a given task, providing a powerful performance metric for object detection.
@@
In Table I, mAP-50 metrics for bounding boxes in the test set are shown. The YOLOv8m model achieved the highest scores, with 0.789 for caruru weed objects, 0.628 for grassy weed objects, and 0.887 for Soy objects. Notably, the YOLOv8 models outperformed the YOLOv5 models across all classes, indicating a significant improvement.
Table II presents the average bounding box precision and recall metrics for all classes across all models. YOLOv5m exhibited the highest average precision at 0.773, while YOLOv5l achieved the best recall at 0.738. Notably, YOLOv8 and YOLOv5 models demonstrated similar results in these metrics.
@@
The segmentation mAP-50 values can be seen in Table III, with YOLOv8m consistently outperforming other models across all classes. Specifically, YOLOv8m achieved a mAP of 0.787 for caruru and 0.696 for grassy weed, and an impressive 0.901 for Soy. Once more, the YOLOv8 models demonstrated superior performance compared to the YOLOv5 models.
@@ VI Analysis of performance on initial and late stages
In this section, the primary focus revolves around the neural network’s role in effectively detecting weed presence. Our analysis will be dedicated to evaluating the performance of the YOLOv8 models on images featuring a coexistence of both weed and soy elements, through box predictions rather then mask predictions. We curated two distinct sets of 25 images, each depicting varying degrees of weed presence. The initial group comprises images captured during the seedling and early stages of plantation, while the second group encapsulates images from the harvest stages. Our objective is to test the YOLOv8 models’ efficacy in accurately identifying weed instances in these diverse agricultural scenarios, and if there is a several difference in performance from early to late stages.
@@
Considering the box mAP50 results in Table V and Table VI, it is evident that the models perform better when both soy and weeds are present in the image. This improved performance might be attributed to the networks producing false positives in the absence of weeds. However, this is not a significant concern given the task at hand, especially since precision remains high in scenarios where only soy is present.
@@
Another noteworthy observation from the tabulated data is the lack of significant disparities in model performance across different stages of plantation. This indicates that the dataset is diverse and that the trained models can effectively handle a variety of agricultural scenarios.
Considering the box precision and recall metrics in Table VII and Table VIII, it is apparent that no single model distinctly outperforms the others. All YOLOv8 models consistently deliver reliable results in detecting both soy and weed in the images. A subtle trend shows a slight improvement in precision and recall metrics in the later stages of plantation, although this enhancement is not particularly significant.
@@
In conclusion, this analysis highlights the consistent performance of the networks across different stages of plantation, demonstrating the dataset’s diversity and the models’ robust generalization capabilities. Additionally, it reveals a significant performance boost in images featuring both soy and weed, aligning with the primary objective of segmenting and detecting these elements. This reaffirms the models’ effectiveness in achieving their intended purpose.
VII Discussion
In this section, we discuss several key points that emerged from the results and analysis.
Firstly, it is evident that the YOLOv8 models consistently outperform the YOLOv5 models. However, it is intriguing that the YOLOv8m model, with significantly fewer parameters, can sometimes outperform the YOLOv8x model. Our best hypothesis is that the dataset, containing only three classes, may lead to overfitting in the more complex YOLOv8x model, resulting in slightly lower performance compared to the smaller YOLOv8m model in certain cases.
Another important observation is the distinct difference in both box and segmentation mAP50 metrics for different classes. The precision for soy is significantly higher than for the two weed classes. This discrepancy likely arises because the dataset was collected in a soy plantation, resulting in a higher prevalence of soy compared to weeds, both in terms of pixel count and presence percentage in images.
It is crucial to clarify that the primary goal of employing these models was not to achieve state-of-the-art results per se, but to validate our dataset. This is why we utilized existing models rather than developing a new one specifically tailored to this task.
VIII Conclusion